OpenStudy (anonymous):

Tell Whether You Would Factor The Polynimoial By Grouping 4w^2+23w+15

OpenStudy (anonymous):

three terms means no I would not use grouping

OpenStudy (mathmate):

Grouping for three terms can be done as follows: 4w^2+23w+15 Find m,n such that m.n=4*15=60, and m+n=23 The obvious choice is m=20,n=3 Rewrite equation as: 4w^2+23w+15 = 4w^2+20w 3w + 15 which factorizes to (w+5)(4w+3)